if a,b,c,d are natural numbers,and a=bc,b=cd,c=da and d=ab then (a+b)(b+c)(c+d)(d+a)=
 

Grade:8

1 Answers

Rakesh mehta
7 years ago
got it.
multiplying all;abcd=1
a/b=bc/cd
a/b=b/d d=ab
a*ab=b^2
a^2=b
By substituting equation ,
d=a^3
c=a^4
b=a^2
a=a
multiplying these all
abcd=a^10
as we have earlier
abcd=1
1=a^10
a=1
b=1
c=1
d=1
solution=
16